# Ear-Keeper: Real-time diagnosis of ear lesions utilizing ultralight-ultrafast ConvNet and large-scale ear endoscopic dataset

### v4

@@ Abstract
Deep learning-based ear disease diagnosis technology has proven effective and affordable. However, due to the lack of ear endoscope datasets with diversity, the practical potential of the deep learning model has not been thoroughly studied. Moreover, existing research failed to achieve a good trade-off between model inference speed and parameter size, rendering models inapplicable in real-world settings. To address these challenges, we constructed the first large-scale ear endoscopic dataset comprising eight types of ear diseases and disease-free samples from two institutions. Inspired by ShuffleNetV2, we proposed Best-EarNet, an ultrafast and ultralight network enabling real-time ear disease diagnosis. Best-EarNet incorporates a novel Local-Global Spatial Feature Fusion Module and multi-scale supervision strategy, which facilitates the model focusing on global-local information within feature maps at various levels. Utilizing transfer learning, the accuracy of Best-EarNet with only 0.77M parameters achieves 95.23% (internal 22,581 images) and 92.14% (external 1,652 images), respectively. In particular, it achieves an average frame per second of 80 on the CPU. From the perspective of model practicality, the proposed Best-EarNet is superior to state-of-the-art backbone models in ear lesion detection tasks. Most importantly, Ear-keeper, an intelligent diagnosis system based Best-EarNet, was developed successfully and deployed on common electronic devices (smartphone, tablet computer and personal computer). In the future, Ear-Keeper has the potential to assist the public and healthcare providers in performing comprehensive scanning and diagnosis of the ear canal in real-time video, thereby promptly detecting ear lesions.
Keywords Ear Diseases  $\cdot$ Deep Learning  $\cdot$ Feature Fusion  $\cdot$ Light-weight Network  $\cdot$ Intelligent System
@@ 3.3 Metrics adopted in this work
[…]

### v5

# Ear-Keeper: A Cross-Platform AI System for Rapid and Accurate Ear Disease Diagnosis

@@ Abstract
Early and accurate detection systems for ear diseases, powered by deep learning, are essential for preventing hearing impairment and improving population health. However, the limited diversity of existing otoendoscopy datasets and the poor balance between diagnostic accuracy, computational efficiency, and model size have hindered the translation of artificial intelligence (AI) algorithms into healthcare applications. In this study, we constructed a large-scale, multi-center otoendoscopy dataset covering eight common ear diseases and healthy cases. Building upon this resource, we developed Best-EarNet, an ultrafast and lightweight deep learning architecture integrating a novel Local–Global Spatial Feature Fusion Module with a multi-scale supervision strategy, enabling real-time and accurate classification of ear conditions. Leveraging transfer learning, Best-EarNet, with a model size of only 2.94 MB, achieved diagnostic accuracies of 95.23% on an internal test set (22,581 images) and 92.14% on an external test set (1,652 images), while requiring only 0.0125 seconds (80 frames per second) to process a single image on a standard CPU. Further subgroup analysis by gender and age showed consistently excellent performance of Best-EarNet across all demographic groups. To enhance clinical interpretability and user trust, we incorporated Grad-CAM-based visualization, highlighting the specific abnormal ear regions contributing to AI predictions. Most importantly, we developed Ear-Keeper, a cross-platform intelligent diagnosis system built upon Best-EarNet, deployable on smartphones, tablets, and personal computers. Ear-Keeper enables public users and healthcare providers to perform comprehensive real-time video–based ear canal screening, supporting early detection and timely intervention of ear diseases.
Keywords Ear Diseases  $\cdot$ Deep Learning  $\cdot$ Health care  $\cdot$ Intelligent System
@@ 3.3 Metrics adopted in this work
$R=\frac{{TP}}{{TP+FN}}{\kern 1.0pt}({\rm{6}}){\kern 1.0pt}{\kern 1.0pt}{\kern 1.0pt}{\kern 1.0pt}{\kern 1.0pt}{\kern 1.0pt}$